Hello,
some of you may remember that 7 years ago I purchased a HP 1707A oscilloscope, which worked for a little bit and then stopped working. I also wrote a topic looking for help for its repair (
https://www.eevblog.com/forum/repair/help-with-hp-1707a-oscilloscope-repair/). Now, it's been almost 5 years since I put my hands on the scope, a lot of things have changed (like getting a EE degree ahah) and I've acquired other gear.
Finding myself stuck in the current lockdown situation, I've decided to start from scratch with A COMPLETE RESTORATION of the oscilloscope. So this thread will be both a journal for the whole trip and also a place where to ask for help (my EE knowledge expanded through the years, but I can't know everything!).
What is wrong with the scope:
- Power supply: in the original post from 7 years ago, the main suspect was the HV section; now I think it's probably a problem in the LV converter board.
- some pot knobs are missing/disintegrated/in bad shape. For sure, one vertical amplitude knob, the trigger knob and the time base selector knob need to be replaced (they are missing).
- internal cabling checking needed: the almost 50 yo heatshrink is slowly melting away.
Available test equipment for the project:
- several multimeters, lab power supply, function generator, one channel 10 MHz scope, one of those 20$ all in one component testers, an ESR meter, an octopus component tester (most of this stuff is DIY).
I also found the service manual for this very scope. Here's the link:
http://hparchive.com/Manuals/HP-1707A-Manual-75Mhz-Scope.pdfIn the following days/weeks I will post updates.